OpenStudy (anonymous):

2. Moving waves can be described either as a function of time or as a function of A. position. B. speed. C. amplitude. D. frequency.

OpenStudy (anonymous):

ik d.is Incorrect

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I already answered to that question

OpenStudy (anonymous):

amplitude?

OpenStudy (anonymous):

ik but d. was rong

OpenStudy (anonymous):

D. frequency. is wrong

OpenStudy (michele_laino):

I think position, since a plane wave can be represented by this formula: \[\Large f\left( {x,t} \right) = A{e^{ - \left( {kx - \omega t} \right)}}\] where x is the position, and t is time

OpenStudy (michele_laino):

oops.. \[\Large f\left( {x,t} \right) = A{e^{ - i\left( {kx - \omega t} \right)}}\]

OpenStudy (anonymous):

20. Using the DMS method to describe an angle, one degree of angle measurement can be divided into how many minutes? A. 100′ B. 90′ C. 360′ D. 60′

OpenStudy (anonymous):

c is Incorrect

OpenStudy (anonymous):

this is the last one

OpenStudy (michele_laino):

sorry what is DMS method?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

here can you help me with this one 19. Trigonometric functions using the ratio of two sides of a triangle can be used to solve _______ triangles. A. acute B. right C. obtuse D. oblique

OpenStudy (michele_laino):

I know DMS method stands for: Degrees Minute Seconds method so the last option

OpenStudy (anonymous):

c is wrong

OpenStudy (anonymous):

i think is d

OpenStudy (michele_laino):

yes it is option D